# Temporal Preference Optimization for Long-Form Video Understanding

### v2

@@ Appendix G Qualitative Analysis Examples
We provide three additional qualitative analysis examples from the Video-MME dataset [15], as illustrated in Figure 10. Example (a) involves an information extraction and optical character recognition (OCR) task, where the question asks for the total number of measurements involved in chip manufacturing. The original LongVA model failed to accurately locate the relevant frame containing the necessary information, resulting in an incorrect response. In contrast, our LongVA-TPO model, enhanced through temporal preference optimization, successfully identified the pertinent frame within the lengthy input and provided the correct answer to the question.

### v3

@@ Appendix F Qualitative Analysis Examples
We provide three additional qualitative analysis examples from the Video-MME dataset [14], as illustrated in Fig. 11. Example (a) involves an information extraction and optical character recognition (OCR) task, where the question asks for the total number of measurements involved in chip manufacturing. The original LongVA model failed to accurately locate the relevant frame containing the necessary information, resulting in an incorrect response. In contrast, our LongVA-TPO model, enhanced through temporal preference optimization, successfully identified the pertinent frame within the lengthy input and provided the correct answer to the question.